OpenStudy (anonymous):

In playing Monopoly, rolling doubles three times in a row sends you to jail. What is the probability of rolling three consecutive doubles? a. 0.167 b. 0.5 c. 0.00463 d. 18.0 @ybarrap

OpenStudy (ybarrap):

Doubles in a single toss can happen 6 ways, there are 6*6 possible ways, so on one toss, there is 6/6^2 probability of a double. For this to happen 3 times in a row is even lower: (6/6^2)^3 = 6^2/6^6 =1/6^3 = .00463